The fact that generally accepted accounting principles allow companies flexibility in choosing between certain allocation methods can make it difficult for a financial analyst to compare periodic performance from firm to firm.

Suppose you were a financial analyst trying to compare the performance of two companies. Company A uses the double-declining-balance depreciation method. Company B uses the straight-line method. You have the following information taken from the 12/31/2021 year-end financial statements for Company B:

Income Statement
Depreciation expense $ 12,500
Balance Sheet
Assets:
Plant and equipment, at cost $ 125,000
Less: Accumulated depreciation (50,000 )
Net $ 75,000


You also determine that all of the assets constituting the plant and equipment of Company B were acquired at the same time, and that all of the $125,000 represents depreciable assets. Also, all of the depreciable assets have the same useful life and residual values are zero.

Required:

1. In order to compare performance with Company A, estimate what B's depreciation expense would have been for 2021 if the double-declining-balance depreciation method had been used by Company B since acquisition of the depreciable assets.

Double-declining balance Year 1 (2018) Year 2 (2019) Year 3 (2020) Year 4 (2021)
2. If Company B decided to switch depreciation methods in 2021 from the straight line to the double-declining-balance method, prepare the 2021 journal entry to record depreciation for the year, assuming no journal entry for depreciation in 2021 has yet been recorded.
Answer #1

ANS:

1. 2018 - $1,25,000*10% = $12,500*2 = $25,000.00

2019 - $1,00,000*10% = $10,000*2 = $20,000.00

2020 - $80,000*10% = $8,000*2 = $16,000.00

2021 - $64,000*10% = $6,400*2 = $12,800.00

2. Accumulated Depreciation A/c Dr. $23,800.00

Plant & Machinery Cr. $23,800.00

(being Additional depreciation for 2018-19-20-21 has been claimed in 31.03.2021)